# Bayesian Signal Matching for Transfer Learning in ERP-Based Brain Computer Interface

### v1

@@ 3.1 Model Reparametrization and Prior Specifications
$\quad\quad\quad\quad\bm{\Sigma}_{0}^{s}Z_{n}+\bm{\Sigma}_{n}^{s}(1-Z_{n}),$
$\quad\quad\quad\quad\bm{R}_{0}^{t}Z_{n}+\bm{R}_{n}^{t}(1-Z_{n})\},\quad n>0,$
where $\bm{B}_{n,1}=(\bm{\beta}_{n,1,1}^{\top},\cdots,\bm{\beta}_{n,1,E}^{\top})$ and $\bm{B}_{0,0}=(\bm{\beta}_{0,0,1}^{\top},\cdots,\bm{\beta}_{0,0,E}^{\top})$.
@@
where $00\leq\gamma_{0}<2,s_{0}>$ 0. In practice, $\gamma_{0}$ and $s_{0}$ are treated as hyper-parameters and selected among a group of values and we do not further differentiate among participants. For participant-specific kernel variance parameters $\psi_{n,1}$ and $\psi_{0,0}$, we assign Log-Normal priors with the a mean zero and a scale one. In addition, we define a function $\phi(\cdot)$ as an eigenfunction of kernel $k$ with eigenvalue $\lambda$ if it satisfies the following integral:
@@
In general, the number of eigenfunctions is infinite, i.e., $\{\phi_{l}(s)\}_{l=1}^{+\infty}$, and we rank them by the descending order of their corresponding eigenvalues $\{\lambda_{l}\}_{l=1}^{+\infty}$. The Mercer’s Theorem suggests that the kernel function $k$ can be decomposed as a weighted summation of normalized eigenfunctions $\phi_{l}(s)$ with positive eigenvalues $\lambda_{l}$ and approximated with a sufficient large $L$ subgroup as follows:
$k(s,s^{\prime})=\sum_{l=1}^{+\infty}\lambda_{l}\phi_{l}(s)\phi_{l}(s^{\prime})\approx\sum_{l=1}^{L}\lambda_{l}\phi_{l}(s)\phi_{l}(s^{\prime})$
Thus, a random function $\bm{f}$ can be represented as a linear combination of the above eigenfunctions and approximated with a sufficient large $L$ subgroup evaluated at $s$ as follows:
$\bm{f}(s)$ $=\begin{pmatrix}f_{1}(s)\\ \vdots\\ f_{e}(s)\\ \vdots\\ f_{E}(s)\\ \end{pmatrix}=\begin{pmatrix}\sum_{l=1}^{\infty}a_{1,l}\phi_{l}(s)\\ \vdots\\ \sum_{l=1}^{\infty}a_{e,l}\phi_{l}(s)\\ \vdots\\ \sum_{l=1}^{\infty}a_{E,l}\phi_{l}(s)\\ \end{pmatrix}\approx\begin{pmatrix}\sum_{l=1}^{L}a_{1,l}\phi_{l}(s)\\ \vdots\\ \sum_{l=1}^{L}a_{e,l}\phi_{l}(s)\\ \vdots\\ \sum_{l=1}^{L}a_{E,l}\phi_{l}(s)\\ \end{pmatrix},$
where $(a_{1,l},\cdots,a_{E,l})^{\top}\sim\mathcal{MVN}(\bm{0},\bm{\Omega}),l>0$. For participant-and-channel-specific noise variance parameters $\sigma_{n,e}^{2}$, we assign Half-Cauchy priors with a mean zero and a scale five. For temporal and spatial parameters $\rho_{n}$ and $\eta_{n}$, we assign discrete uniform priors among a grid of candidate values between 0 and 1. Finally, the prior specifications are summarized as follows:
@@
$\psi_{n,1},\psi_{0,0}\overset{\mathrm{indep.}}{\sim}\mathcal{LN}(0,1),$
$\sigma_{n,e}\sim\mathcal{HC}(0,5.0),\quad\rho_{n}\sim\mathcal{U}(0,1),\quad\eta_{n}\sim\mathcal{U}(0,1),$
$\mathrm{for}\quad$ $n=0,\cdots,N,\quad e=1,\cdots,E,$
@@ 3.2 Markov Chain Monte Carlo
where $\bm{A}_{n,1}$ and $\bm{A}_{0,0}$ follow matrix normal prior distributions with closed posterior forms (Zhang et al.,, 2021), and $\bm{\Psi}_{1}^{1:L_{1}},\bm{\Psi}_{0}^{1:L_{0}},\bm{V}_{\bm{\Psi}_{1}}^{1:L_{1}},$ and $\bm{V}_{\bm{\Psi}_{0}}^{1:L_{0}}$ are the first $L_{1}$ rows of $\bm{\Psi}_{1}$, the first $L_{0}$ rows of $\bm{\Psi}_{0}$, the diagonal matrix of the first $L_{1}$ eigenvalues associated with $\bm{\Psi}_{1}$, and the diagonal matrix of the first $L_{0}$ eigenvalues associated with $\bm{\Psi}_{0}$, respectively. In practice, $L_{1}$ and $L_{0}$ are determined by the minimum integer of which cumulative sum of eigenvalues divided by the total sum of eigenvalues is over 95% after reordering them by the descending order. For the remaining parameters, we apply Metropolis-Hastings algorithm to draw posterior samples. For the convergence check, we run multiple chains with different seed values and evaluate the convergence by Gelman-Rubin statistic (Gelman and Rubin,, 1992). Details of MCMC implementation can be found in the supplementary material.
Figure 2: An illustration of the BSM-Mixture algorithm. First, we pre-defined a threshold $\delta_{Z}\in(0,1)$. Second, we fitted BSM model and obtained posterior samples of $\{Z_{n}\}$ to compute $\mathbb{E}(Z_{n}\mid\bm{\Theta})$. If none of the posterior expectations exceeded $\delta_{Z}$, we applied BSM-Reference, otherwise, we applied BSM model for prediction. Note that both BSM-Reference and BSM models can be fitted simultaneously, and it will not have extra computational burden to the total calibration effort.
@@ 3.3 Hybrid Model Selection Strategy: BSM-Mixture
Since our primary goal was to use small amount of training data from the new participant for calibration data, it might be challenging for our BSM model to identify source participants when no source participants were available for matching. In practice, we proposed the hybrid model selection strategy, namely BSM-Mixture, such that if no source participants were selected satisfying certain threshold, we simply used BSM-Reference for prediction. We defined the hybrid model selection strategy as follows: First, we pre-defined a threshold $\delta_{Z}\in(0,1)$. Second, we fitted BSM model and obtained the posterior samples of $\{Z_{n}\}$ to compute $\mathbb{E}(Z_{n}\mid\bm{\Theta})$. If none of $\mathbb{E}(Z_{n}\mid\bm{\Theta})$ exceeded $\delta_{Z}$, we applied BSM-Reference, otherwise, we applied BSM model for prediction. Note that BSM-Reference and BSM model could be fitted simultaneously, it did not add much computational burden to the total calibration effort. Figure 2 provides an illustration of the hybrid model selection strategy on BSM-Mixture.
@@ 4 Simulation Studies
We conducted extensive simulation studies to demonstrate the advantage of our proposed framework. We showed the results of the multi-channel simulation scenario in this section and moved the results of the single-channel scenario to the supplemental material.
@@
Setup  We considered the scenario with $N=7$ and $K=3$. We designed three groups of parameters including two-dimensional pre-specified mean ERP functions and variance nuisance parameters. The shape and magnitude were based on real participants from the database (Thompson et al.,, 2014). The simulated EEG signals were generated with a response window of 35 time points per channel, i.e., $T_{0}=35$. For simulated data generated from group 0, we created a typical P300 pattern for the first channel where the target ERP function reached its positive peak around 10th time point post stimulus and reversed the sign for both ERP functions for the second channel; for simulated data generated from group 1, we created the same target ERP function as group 0 for the first channel and reduced the peak magnitude for the second channel; for simulated data generated from group 2, we reduced the peak magnitude for the first channel and created the same ERP functions as group 1 for the second channel. We considered an autoregressive temporal structure of order 1 (i.e., AR(1)), compound symmetry spatial structure, and channel-specific variances for background noises, where the true parameters for groups 0, 1, and 2 were $(\rho_{0}=0.7,\eta_{0}=0.6,\sigma_{0,1}=8.0,\sigma_{0,2}=8.0)$, $(\rho_{1}=0.7,\eta_{1}=0.4,\sigma_{1,1}=8.0,\sigma_{1,2}=6.0)$, and $(\rho_{2}=0.5,\eta_{2}=0.4,\sigma_{2,1}=2.0,\sigma_{2,2}=2.0)$, respectively. We still designed two cases for this scenario, a case without matched data among source participants (Case 1) and a case with matched data (Case 2). For Case 1, the group labels for source participants 1-6 were 1, 1, 1, 2, 2, and 2, respectively. For Case 2, the group labels for source participants 1-6 were 0, 0, 1, 1, 2, and 2, respectively. We performed 100 replications for each case. Within each replication, we assumed that each participant was spelling three characters “TTT” with ten sequence replications per character for training and generated additional testing data of the same size as the testing data of the single-channel scenario.
@@
Settings and Diagnostics  All simulated datasets were fitted with equation (6). We also applied two covariance kernels $\kappa_{1}$ and $\kappa_{0}$ for target and non-target ERP functions, respectively. The two kernels were characterized by $\gamma$-exponential kernels with length-scale and gamma hyper-parameters as $(0.2,1.2)$ and $(0.3,1.2)$, respectively. The pre-specified threshold $\delta_{Z}$ was $0.5$. We ran the MCMC with three chains, with each chain containing 3000 burn-ins and 300 MCMC samples. The Gelman-Rubin statistics were evaluated.
Table 1: The means and standard deviations of testing sequence sizes to achieve a 95% testing accuracy given certain training sequence sizes for BSM-Mixture, BSM, MDWM, BSM-Reference, and swLDA across 100 replications. For Case 1, the testing sequence sizes of BSM-Mixture were smaller or close to MDWM and swLDA although the testing sequence sizes of BSM were consistently larger than those of BSM-Reference, suggesting that the strategy of BSM-Mixture favored BSM-Reference when no one was selected; For Case 2, the testing sequence sizes of BSM-Mixture were consistently smaller than MDWM and swLDA, and the strategy of BSM-Mixture favored BSM when at least one source participant was selected.
Train Test Sequence Size
Seq Size BSM-Mixture BSM MDWM BSM-Reference swLDA
Case 1 2 8.4, 1.7 9.8, 0.7 7.9, 1.6 8.2, 1.7 10.0, 0.3
3 7.6, 1.8 9.6, 1.1 7.8, 1.6 7.6, 1.8 9.8, 0.6
4 7.2, 1.7 9.3, 1.1 7.8, 1.4 7.1, 1.7 9.6, 0.9
Case 2 2 7.3, 1.6 7.3, 1.7 7.9, 1.6 8.2, 1.7 10.0, 0.3
3 7.2, 1.8 7.2, 1.8 7.9, 1.6 7.6, 1.8 9.8, 0.6
4 6.8, 1.7 6.8, 1.7 7.8, 1.5 7.1, 1.7 9.6, 0.9
Table 2: The mean and standard deviations of training sequence size to achieve a 95% testing accuracy given certain testing sequence sizes for BSM-Mixture, BSM, MDWM, BSM-Reference, and swLDA across 100 iterations. For Case 1, the training sequence sizes of BSM-Mixture were consistently smaller than MDWM and swLDA although BSM had larger training sequence sizes than BSM-Reference, suggesting that the strategy of BSM-Mixture favored BSM-Reference when no one was selected; For Case 2, the training sequence sizes of BSM-Mixture are consistently smaller than MDWM and swLDA, and the strategy of BSM-Mixture favored BSM when at least one source participant was selected.
Test Train Sequence Size
Seq Size BSM-Mixture BSM MDWM BSM-Reference swLDA
Case 1 5 8.0, 3.0 9.2, 1.7 9.0, 2.5 8.0, 3.1 9.7, 1.2
6 6.3, 3.4 8.3, 2.2 7.5, 3.5 6.3, 3.5 9.2, 1.6
7 4.9, 3.4 7.1, 2.9 5.2, 3.9 4.9, 3.5 8.7, 2.1
Case 2 5 7.5, 3.5 7.4, 3.5 8.9, 2.6 8.0, 3.1 9.7, 1.2
6 5.7, 4.1 5.8, 4.1 7.5, 3.4 6.3, 3.5 9.2, 1.6
7 3.8, 3.7 3.9, 3.7 5.3, 4.0 4.9, 3.5 8.7, 2.1
Criteria  We evaluated our method by matching and prediction. For matching, we reported the proportion that each source participant was matched to the new participant and produced the ERP function estimates with 95% credible bands with respect to the training sequence size. For prediction, we reported the character-level prediction accuracy of the testing data, using Bayesian signal matching with source participants’ data (BSM), Bayesian generative methods with the new participant’s data only (BSM-Reference), the hybrid model selection strategy (BSM-Mixture), an existing classification method with data borrowing using Riemannian Geometry (MWDM), and swLDA with new participant’s data only (swLDA). We did not include SMGP for comparison in the simulation studies because SMGP assumed that data were generated under sequences of stimuli, while our data generative mechanism was based on each stimulus. We did not include other conventional ML methods because we assume that the prediction accuracy of swLDA was representative of those types of methods.
@@
Table 3: Upper Panel: A summary of means and standard deviations of $Z_{n}=1$ for Case 1 across 100 replications. Lower Panel: A summary of means and standard deviations of $Z_{n}=1$ for Case 2 across 100 replications. The results were produced by BSM. The numerical values were multiplied by 100 for convenience and were reported with respect to the training sequence replications. For Case 1, all the values quickly decreased below 10.0 with 3 training sequence replications; for Case 2, values associated with Participants 1-2 were above 0.75 across all training sequence replications and values associated with Participants 3-6 were below 10.0 within 4 sequence replications. The results indicated that our method correctly performed the participant selection.
@@
Matching Results  The upper and lower panels of Table 3 show the means and standard deviations of $Z_{n}=1$ across 100 replications for cases without and with matched data, respectively, by BSM. The numerical values were multiplied by 100 for convenient reading and were reported with respect to the training sequence size of the new participant. For Case 1, all the values quickly dropped below 10.0 with 3 training sequence replications; for Case 2, values associated with Participants 1-2 were above 75.0 across all the training sequences and values associated with Participants 3-6 quickly dropped below 10.0 within 4 sequence replications. The results indicated that our method correctly performed the participant selection.
@@
Figure 3: Means and standard errors of testing prediction accuracy by BSM-Mixture, BSM, BSM-Reference, MDWM, and swLDA of Case 1. The upper, middle, and lower panels show prediction accuracy with the first 2, 3, and 4 training sequence replications, respectively.
Figure 4: Means and standard errors of testing prediction accuracy by BSM-Mixture, BSM, BSM-Reference, MDWM, and swLDA of Case 2. The upper, middle, and lower panels show prediction accuracy with the first 2, 3, and 4 training sequence replications, respectively.
@@
Prediction Results  Figure 3 shows the means and standard errors of testing prediction accuracy of Case 1 (No Matched) by BSM-Mixture, BSM, BSM-Reference, MDWM, and swLDA. The testing prediction accuracy is further stratified by training the first 2, 3, and 4 sequences of the new participant. Within the same row, BSM-Mixture and BSM-Reference had similar performances to MDWM. The poor performance of BSM was due to the uncertainty among participant selection in presence of no source participant and small training sample size. Therefore, the hybrid model selection strategy solved this issue. Within the same column, all four methods themselves performed better as training data of the new participant increased.
Figure 4 shows the means and standard errors of testing prediction accuracy of Case 2 (Matched) by BSM-Mixture, BSM, BSM-Reference, MWDM, and swLDA. The testing prediction accuracy is also stratified by training the first 2, 3, and 4 training sequences of the new participant. Within the same row, BSM-Mixture performed the best, followed by MDWM, suggesting that our method successfully identified the matched participants and leveraged their information correctly. Within the same column, all four methods themselves performed better as training data of the new participant increased.

### v2

@@ 3.1 Model Reparametrization and Prior Specifications
$\quad\quad\quad\quad\bm{\Sigma}_{0}^{s}Z_{n}+\bm{\Sigma}_{n}^{s}(1-Z_{n}),\bm{R}_{0}^{t}Z_{n}+\bm{R}_{n}^{t}(1-Z_{n})\},$
where $\bm{B}_{n,1}=(\bm{\beta}_{n,1,1}^{\top},\cdots,\bm{\beta}_{n,1,E}^{\top})(n>0)$ and $\bm{B}_{0,0}=(\bm{\beta}_{0,0,1}^{\top},\cdots,\bm{\beta}_{0,0,E}^{\top})$.
@@
$\bm{f}(s)$ $=\begin{pmatrix}f_{1}(s)\\ \vdots\\ f_{E}(s)\\ \end{pmatrix}=\begin{pmatrix}\sum_{l=1}^{\infty}a_{1,l}\phi_{l}(s)\\ \vdots\\ \sum_{l=1}^{\infty}a_{E,l}\phi_{l}(s)\\ \end{pmatrix}\approx\begin{pmatrix}\sum_{l=1}^{L}a_{1,l}\phi_{l}(s)\\ \vdots\\ \sum_{l=1}^{L}a_{E,l}\phi_{l}(s)\\ \end{pmatrix},$
where $(a_{1,l},\cdots,a_{E,l})^{\top}\sim\mathcal{MVN}(\bm{0},\bm{\Omega}),l>0$. For participant-and-channel-specific noise variance parameters $\sigma_{n,e}^{2}$, we assign Half-Cauchy priors with a mean zero and a scale five. For temporal and spatial parameters $\rho_{n}$ and $\eta_{n}$, we assign discrete uniform priors among a grid of candidate values between 0 and 1. For $n=0,\cdots,N,e=1,\cdots,E$,
@@
$\psi_{n,1},\psi_{0,0}\overset{\mathrm{indep.}}{\sim}\mathcal{LN}(0,1),\quad\sigma_{n,e}\sim\mathcal{HC}(0,5.0),\quad\rho_{n}\sim\mathcal{U}(0,1),\quad\eta_{n}\sim\mathcal{U}(0,1),$
@@ 3.2 Markov Chain Monte Carlo
where $\bm{A}_{n,1}$ and $\bm{A}_{0,0}$ follow matrix normal prior distributions with closed posterior forms (Zhang et al.,, 2021), and $\bm{\Psi}_{1}^{1:L_{1}},\bm{\Psi}_{0}^{1:L_{0}},\bm{V}_{\bm{\Psi}_{1}}^{1:L_{1}},$ and $\bm{V}_{\bm{\Psi}_{0}}^{1:L_{0}}$ are the first $L_{1}$ rows of $\bm{\Psi}_{1}$, the first $L_{0}$ rows of $\bm{\Psi}_{0}$, the diagonal matrix of the first $L_{1}$ eigenvalues associated with $\bm{\Psi}_{1}$, and the diagonal matrix of the first $L_{0}$ eigenvalues associated with $\bm{\Psi}_{0}$, respectively. In practice, $L_{1}$ and $L_{0}$ are determined by the minimum integer of which cumulative sum of eigenvalues divided by the total sum of eigenvalues is over 95% after reordering them by the descending order. For the remaining parameters, we apply Metropolis-Hastings algorithm to draw posterior samples. For the convergence check, we run multiple chains with different seed values and evaluate the convergence by Gelman-Rubin statistic (Gelman and Rubin,, 1992). In addition, we examine the chain-specific posterior means of $\{Z_{n}\}$ as a post-hoc way to demonstrate the convergence. Details of MCMC implementations can be found in the supplementary material.
Figure 2: An illustration of the BSM-Mixture algorithm. First, we pre-define a threshold $\delta_{Z}\in(0,1)$. Second, we fit BSM model and obtain posterior samples of $\{Z_{n}\}$ to compute $\mathbb{E}(Z_{n}\mid\bm{\Theta})$. If none of the posterior expectations exceed $\delta_{Z}$, we apply BSM-Reference, otherwise, we apply BSM model for prediction. Note that both BSM-Reference and BSM models can be fitted simultaneously, and it will not have extra computational burden to the total calibration effort.
@@ 3.3 Hybrid Model Selection Strategy: BSM-Mixture
Since our primary goal is to use small amount of training data from the new participant for calibration data, it may be challenging for our BSM model to identify source participants when no source participants are available for matching. In practice, we propose the hybrid model selection strategy, namely BSM-Mixture, such that if no source participants were selected satisfying certain threshold, we would use BSM-Reference for prediction. We define the hybrid model selection strategy as follows: First, we pre-define a threshold $\delta_{Z}\in(0,1)$. Second, we fit BSM model and obtain the posterior samples of $\{Z_{n}\}$ to compute $\mathbb{E}(Z_{n}\mid\bm{\Theta})$. If none of $\mathbb{E}(Z_{n}\mid\bm{\Theta})$ exceed $\delta_{Z}$, we apply BSM-Reference, otherwise, we apply BSM model for prediction. Note that BSM-Reference and BSM model can be fitted simultaneously, it does not add much computational burden to the total calibration effort. In addition, since the current data have been collected from the single lab source, we do not expect to see significant impact of machine effect on the parameter estimationa and prediction accuracy. Figure 2 provides an illustration of the hybrid model selection strategy on BSM-Mixture.
@@ 4 Simulation Studies
We conduct extensive simulation studies to demonstrate the advantage of our proposed framework. We show the results of two simulation scenarios: a naive multi-channel setting and a real data based setting.
4.1 Naive Multi-channel Simulation Scenario
@@
Setup  We consider the scenario with $N=7$ and $K=3$. The simulated EEG signals are generated from multivariate normal distributions with a response window of 35 time points per channel, i.e., $T_{0}=35$. We design three groups of parameters including two-dimensional pre-specify mean ERP functions and variance nuisance parameters, where the shape and magnitude parameters are based on real participants from the database (Thompson et al.,, 2014). For simulated data generated from group 0, we create a typical P300 pattern for the first channel where the target ERP function reach its positive peak around 10th time point post stimulus and reverse the sign for both ERP functions for the second channel; for simulated data generated from group 1, we create the same target ERP function as group 0 for the first channel and reduce the peak magnitude for the second channel; for simulated data generated from group 2, we reduce the peak magnitude for the first channel and create the same ERP functions as group 1 for the second channel. We consider an autoregressive temporal structure of order 1 (i.e., AR(1)), compound symmetry spatial structure, and channel-specific variances for background noises, where the true parameters for groups 0, 1, and 2 are $(\rho_{0}=0.7,\eta_{0}=0.6,\sigma_{0,1}=8.0,\sigma_{0,2}=8.0)$, $(\rho_{1}=0.7,\eta_{1}=0.4,\sigma_{1,1}=8.0,\sigma_{1,2}=6.0)$, and $(\rho_{2}=0.5,\eta_{2}=0.4,\sigma_{2,1}=2.0,\sigma_{2,2}=2.0)$, respectively. We design two cases for this scenario, a case without matched data among source participants (Case 1) and a case with matched data (Case 2). For Case 1, the group labels for source participants 1-6 are 1, 1, 1, 2, 2, and 2, respectively. For Case 2, the group labels for source participants 1-6 were 0, 0, 1, 1, 2, and 2, respectively. We perform 100 replications for each case. Within each replication, we assume that each participant is spelling the character “T” three times with ten sequence replications per character for training and generate additional testing data of the same size as the testing data of the single-channel scenario.
@@
Settings and Diagnostics  All simulated datasets are fitted with equation (5). We also apply two covariance kernels $\kappa_{1}$ and $\kappa_{0}$ for target and non-target ERP functions, respectively. The two kernels are characterized by $\gamma$-exponential kernels with length-scale and gamma hyper-parameters as $(0.2,1.2)$ and $(0.3,1.2)$, respectively. The pre-specified threshold $\delta_{Z}$ is $0.5$. We run the MCMC with three chains, with each chain containing 5,000 burn-ins and 3,000 MCMC samples.
@@
Table 1: The upper and lower panels show summary of means and standard deviations of $Z_{n}=1$ for Cases 1 and 2 across 100 replications by BSM, respectively. The numerical values are multiplied by 100 for convenience and were reported with respect to the training sequence replications.
@@
Criteria  We evaluate our method by matching and prediction. For matching, we report the proportion that each source participant is matched to the new participant and produce the ERP function estimates with 95% credible bands with respect to the training sequence size. For prediction, we report the character-level prediction accuracy of the testing data, using Bayesian signal matching with source participants’ data (BSM), Bayesian generative methods with the new participant’s data only (BSM-Reference), the hybrid model selection strategy (BSM-Mixture), an existing classification method with data borrowing using Riemannian Geometry (MWDM), and swLDA with new participant’s data only (swLDA). We do not include SMGP for comparison in the simulation studies because SMGP assumes that data are generated under sequences of stimuli, while our data generative mechanism is based on stimuli. We does not include other conventional ML methods because the prediction accuracy of swLDA is representative of those types of methods.
@@
Matching Results  The upper and lower panels of Table 1 show the means and standard deviations of $Z_{n}=1$ across 100 replications for cases without and with matched data, respectively, by BSM. The numerical values are multiplied by 100 for convenient reading and are reported with respect to the training sequence size of the new participant. For Case 1, all the values quickly drop below 10.0 with 3 training sequence replications; for Case 2, values associated with Participants 1-2 are above 75.0 across all the training sequences and values associated with Participants 3-6 quickly drop below 10.0 within 4 sequence replications. The results indicate that our method correctly perform the participant selection.
@@
Figure 3: The left and right panels show the means and standard errors of testing prediction accuracy of Cases 1 and 2 by BSM-Mixture, BSM, BSM-Reference, MDWM, and swLDA across 100 replications, respectively. The upper, middle, and lower panels show prediction accuracy with the first 2, 3, and 4 training sequence replications, respectively.
@@
Prediction Results  The left and right panels of Figure 3 show the means and standard errors of testing prediction accuracy of Case 1 (No Matched) and Case 2 (Matched) by BSM-Mixture, BSM, BSM-Reference, MDWM, and swLDA, respectively. The testing prediction accuracy is further stratified by training the first 2, 3, and 4 sequences of the new participant. For Case 1, within the same row, BSM-Mixture and BSM-Reference have similar performances to MDWM. The poor performance of BSM is due to the uncertainty among participant selection in presence of no source participant and small training sample size. Therefore, the hybrid model selection strategy solves this issue. For Case 2, within the same row, BSM-Mixture performs the best, followed by MDWM, suggesting that our method successfully identifies the matched participants and leverages their information correctly. Within the same column, four methods benefit from an increasing number of training data from the new participant.
4.2 Real Data Based Simulation Scenario
We design a real-data based multi-channel simulation scenario. True parameters and sample sizes of training sets are copied from the real data analysis. We follow the tradition that participant 0 refers to the new participant, while participants 1 to 23 refer to the source participants. The participant-specific mean ERP functions and variance nuisance parameters are obtained from posterior means of each participant in the real data analysis. We multiply the variance estimates by 2 in this simulation scenario. The simulated EEG signals are also generated from multivariate normal distributions with a response window of 25 time points per channel, i.e., $T_{0}=25$. We perform 100 replications for this scenario. Within each replication, we assume that each participant is spelling the character “T” eight times with ten sequence replications per character for training and generate additional testing data of the same size as the testing data of the naive multi-channel scenario. All simulated datasets are fitted with equation (5). We used the same kernels, hyper-parameters, and the threshold $\delta_{Z}$ as the real data analysis. We run the MCMC with three chains, with each chain containing 5,000 burn-ins and 3,000 MCMC samples. Both matching and prediction results are reported, and we follow the same criteria as mentioned in Section 4.1.
Figure 4: A graphic of the testing prediction accuracy. The upper, middle, and lower panel corresponds to the prediction accuracy with the first 3, 4, and 5 sequences of training data, respectively. All results are evaluated across 100 replications.
ResultsFor matching, we only observe that participant 1 has a posterior mean of being merged around 0.2, while the posterior means of other source participants are below 0.1. The finding is consistent with the matching from the real data analysis given the same participant ordering and supports our argument that we typically expect one or two matching based on BSM classifier. Details of selection probability can be found in Section S4.2 in the supplementary. For prediction, Figure 4 shows the testing prediction accuracy by the same five methods across 100 replications. The upper, middle, and lower panel are the prediction accuracy by training the first 2, 3, and 4 training sequences of the new participant. Overall, we observe that BSM-Mixture performs better than MDWM and swLDA. The prediction curves of BSM-Mixture and BSM are almost identical, suggesting that the algorithm identifies at least one source participant with the posterior mean of being merged exceeds the threshold $\delta_{Z}=0.1$, which is also consistent with the matching result. The prediction accuracy of BSM is slightly higher and narrower than that of BSM-Reference, possibly because the borrowing information may be limited when only one participant has reached the selection threshold. Finally, when we increase the sample size of training data, the prediction accuracy improves accordingly.